TERAHERTZ SPECTROSCOPY OF FREE RADICALS

\underline{G. WINNEWISSER}, I. Physikalisches Institut, Universität zu Köln, D-50913 Cologne.

The first molecular species detected in space were the radicals CH, CH^{+}, and CN. Free radicals and ions constitute a sizeable fraction of the detected interstellar and circumstellar molecules. Recent detailed mapping of the interstellar distribution of CN reveal the different interstellar distribution between open shell and closed shell molecules. High resolution, broadband scanning spectroscopy with microwave accuracy has been extended into the terahertz region by frequency and phase stabilization of continuously tunable Backward Wave Oscillators (BWOs) at frequencies up to 2 THz. In combination with frequency multiplication techniques and alternatively laser side band techniques, this BWO based method allows broadband scanning spectroscopy to be extended to 3 THz. In case the FIR-laser is frequency and phase stabilized to an atomic clock, microwave accuracy is achieved. With this Cologne terahertz spectrometer the rotational spectra of a large number of closed- and open-shell molecules have been measured in a variety of different discharges cooled to liquid nitrogen temperature. Amongst these are many astrophysically important light hydrides and the fine structure transitions ^{3}P_{2} – ^{3}P_{1} of atomic carbon, ^{12}C and ^{13}C, at 809 341.970(50) MHz and 809 493.7(2) MHz respectively. The latter frequency refers to the F=2.5←1.5 hyperfine transitions. Larger carbon chain radicals e.g. C_{4}, C_{9}, C_{11} have been detected by tunable diode laser spectroscopy using laser ablation techniques. We have observed many astrophysically relevant molecular free radicals in the frequency region up to 1.3 THz. Amongst these are the important light hydride radicals CH, CD; NH, ND; PH, PD; SH, SD; NHH, NHD, NDD; CCH, CCD. In addition, many heavier free radicals have been measured such as CF, CN, CP, NO, OO, PS; SSH and SSD together with their heavier isotopic variants. The detailed study of the higher rotational transitions of the heavier radicals did reveal the gradual change from one Hund’s coupling scheme to the other. Details will be discussed. Some of these new laboratory data have led to interstellar detections, others are pending. Some of the stable species such as CO, NH_{3}, HCl, HCN, and others have been investigated by saturation spectroscopy, resulting in line widths of < 40 kHz, well below the Doppler line widths, and in favourable cases with line center determinations of < 1 kHz at frequencies near 1 THz. With an intracavity millimeter wave spectrometer, the Orotron spectrometer, we have detected in the laboratory the astrophysically relevant molecular van der Waals complexes CO-HH, CO-He, and CO-CO.